OpenStudy (anonymous):

If f(x) = x − 2 and g(x) = −2x − 3, find (f • g)(6)

OpenStudy (anonymous):

If f(x) = −x − 1 and g(x) = 4x − 1, find f[g(−2)]

OpenStudy (anonymous):

1) g(6) g(x)----> -2x -3 = -2(6) -3 = -12 - 3 = -15 for the gf(x) i'm not sure....and i don't wanna give you a wrong answer. :/

OpenStudy (anonymous):

what about for this one?

OpenStudy (anonymous):

If f(x) = −x − 1 and g(x) = 4x − 1, find f[g(−2)]

OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = x − 2, find f[g(−6)].

OpenStudy (anonymous):

first one g(2)= 4x-1 = 4(2) -1 = 8-1 = 7 i will give you another example i did in school: Given that f:x---> 3-x g:x---> \[\frac{ x+2 }{ x-5}\] 1) Calculate g(2) 2) State a value of x for which g(x) is not defined. 3) Derive an expression for gf (x) 4) Calculate the value of \[ f{-1}\] (4) Solution: 1) g(x)----> \[\frac{ x+2 }{ x-5}\] \[\frac{ 2+2 }{ 2-5 }\] g(2)= \[\frac{ 4 }{ -3 }\] or -1.33.... 2) g(x)---> \[\frac{ x+2 }{ x-5 }\] \[x-5\neq0\] \[x \neq5\] 3) g(x): \[\frac{ 1(3-x) +2 }{ 1(3-x)-5 }\] \[\frac{ 3-x+2 }{ 3-x-5 }\] =\[\frac{ 3+2-x }{ 3-x-5}\] = \[\frac{ 5-x }{ -2-x}\] 4) \[f{-1}\] (4) f(x)-----> 3-x y= 3-x 5)Interchange x for y x= 3-y y=3-x therefore, \[f{-1}\] (x)= 3-x \[f{-1}\] (4): \[f{-1}\] (x) = 3-x = 3-4 \[f{-1}\] (4) = -1 I hope i didn't confuse you!

OpenStudy (anonymous):

thanks helped a lot

OpenStudy (anonymous):

your welcome :) try writing it down then reading it because it's confusing with all the spaces!